OpenStudy (anonymous):

Solve: 3^4x-7=4^2x+3

myininaya (myininaya):

I think you mean to say 3^(4x-7)=4^(2x+3) correct me if i'm wrong...

OpenStudy (anonymous):

Yes, that is what I meant to ask.

myininaya (myininaya):

So take ln( ) of both sides! :)

myininaya (myininaya):

Then you will be able to bring the exponents down using ln(x^r)=rln(x) this property!

myininaya (myininaya):

\[(4x-7) \ln(3)=(2x+3) \ln(4)\] Now distribute and ... Put every term with x on one side and every term without x on the other side

myininaya (myininaya):

\[4 \ln(3) x -7 \ln(3)=2 \ln(4) x+3 \ln(4)\] this is the part where you do the: "Put every term with x on one side and every term without x on the other side"

myininaya (myininaya):

Would you like to try to finish this? You can use the draw tool if you want to show me how to finish this and I can fix where you went wrong.

myininaya (myininaya):

or go wrong*

OpenStudy (anonymous):

I'm having a bit of trouble with the "Put every term with x on one side and every term without x on the other side." What exactly does this mean?

myininaya (myininaya):

how do you solve 2x+1=3x-1 2x-3x=-1-1 see i put every term with x on one side and every term without x on the other

OpenStudy (anonymous):

Oh, I get that. But what do you do once you move the x's onto the one side and the non x's onto the other?

myininaya (myininaya):

going from the example I just wrote x(2-3)=-1-1 Then divide both sides by (2-3) x=(-1-1)/(2-3)

myininaya (myininaya):

in others word you are going to factor out that x and you will have x(junk)=something

myininaya (myininaya):

\[4 \ln(3) x -7 \ln(3)=2 \ln(4) x+3 \ln(4) \] So here is where I left you...

myininaya (myininaya):

so one possible way to go about this is to subtract 2 ln(4) x on both sides then add 7 ln(3) on both sides

OpenStudy (anonymous):

Ohhh! I got it! The answer is 7.305... Thanks soooo much for your help! :)

myininaya (myininaya):

Awesome! I'm glad you got it.